# Leveraging the true depth of LLMs

### v2

@@ 6 Limitations
The effectiveness of our approach exhibits notable variations across model scales. Smaller models show reduced benefits, likely due to their less sparse activation patterns and more tightly coupled layer dependencies. This degradation becomes more pronounced as the LP sequence increases, suggesting a practical upper limit to the number of layer pairs that can be effectively parallelized.
Finetuning. While some performance loss can be mitigated through fine-tuning, we were unable to fully recover the baseline model’s performance levels. This suggests fundamental trade-offs between computational efficiency and model capability that cannot be entirely eliminated through optimization.
Determining the ’true’ effective depth—the optimal configuration of parallel layer pairs—remains an open challenge as there is no theoretical framework for predicting the optimal grouping strategy.
@@ 7 Conclusion
In this work, we presented Layer Parallelism, a novel approach that exploits independence patterns between transformer layers to optimize LLM inference. By restructuring the computational graph to enable parallel execution of consecutive layer pairs through tensor parallelism, we achieved substantial speed improvements without model retraining. We show that we can recover ICL accuracy by fine-tuning the parallelized models.
@@
Impact Statement
This paper presents work whose goal is to improve inference serving of LLMs by trading accuracy. Well-intentioned users could find themselves receiving incorrect predictions that could lead to undesirable outcomes.
@@ Acknowledgments
We thank Benjamin Rio for his insightful conversations and valuable input during the preparation of this work. His perspectives influenced several key aspects of the presentation.
Appendix A Ablation: Tokens per second
@@
Appendix B Generalization to multiple GPUs
@@
Figure 9: Layer Parallelism in the case of parallelizing two layers over four accelerators. The stacked layers represent the tensor parallelism, and the colors indicate the processing of different previously contiguous layers. $Q,K,V\text{ and }att\in\mathbb{R}^{T\times\frac{2D}{g}}$, where $D$ is the feature dimension and $g$ is the total number of accelerators.
@@
Appendix C Acceleration source
@@
Figure 10 illustrates flame graphs comparing two consecutive Llama 3.2 3B decoder layers using vanilla tensor parallelism and our LP approach. The profiling data summarized in Table 3 reveals that the primary source of acceleration in our LP method stems from reducing the total number of all reduce synchronization operations across GPUs. Specifically, the vanilla tensor parallel approach performs synchronization at every decoder layer, resulting in higher cumulative synchronization overhead. In contrast, our Layer Parallelism implementation runs pairs of layers simultaneously, effectively halving the number of synchronization points. This reduction in synchronization leads to a significant drop in synchronization time from 100.8ms to 50.7ms, directly contributing to the observed improvement in inference speed.

### v3

@@ 6 Limitations
The effectiveness of our approach exhibits notable variations across model scales. Smaller models show reduced benefits, likely due to their less sparse activation patterns and more tightly coupled layer dependencies. This degradation becomes more pronounced as the LP sequence length increases, suggesting a practical upper limit to the number of layer pairs that can be effectively parallelized.
Regarding the fine-tuning, while some performance loss can be mitigated, we were unable to fully recover the baseline model’s performance levels. This suggests fundamental trade-offs between computational efficiency and model capability that cannot be entirely eliminated through optimization, or that more involved fine-tuning strategies might be required.
Moreover, determining the ’true’ effective depth—the optimal configuration of parallel layer pairs—remains an open challenge as there is no theoretical framework for predicting the optimal grouping strategy.
@@ 7 Conclusion
In this work, we presented Layer Parallelism, a novel approach that exploits independence patterns between transformer layers to optimize LLM inference. By restructuring the computational graph to enable parallel execution of consecutive layer pairs through tensor parallelism, we achieved substantial speed improvements without model retraining. Our method reduced the effective depth of Llama 2 7B by 21% while maintaining 98% of the original performance (without fine-tuning), yielding up to a 1.29x improvement in inference speed for single-token generation with long sequences. Moreover, we show that we can recover some of the lost accuracy through naive fine-tuning.
@@ Acknowledgments
We would like to thank Benjamin Rio for the thoughtful discussions and help running some of the performance experiments. Ramón’s research is supported by META.
Appendix A Theoretical Analysis of Layer Parallelism
This section provides a theoretical justification for Layer Parallelism (LP), analyzing the approximation error introduced by the LP computational graph and connecting it to the empirical observations in the main paper.
A.1 Sequential vs. Layer-Parallel Computation
Consider two consecutive transformer decoder layers $\ell_{k}$ and $\ell_{k+1}$ in a pre-norm architecture. Let $A_{k}(\cdot)$ denote the attention residual and $F_{k}(\cdot)$ the feed-forward residual of layer $k$.
Exact sequential computation.
The standard forward pass computes:
$u_{k}$ $=x+A_{k}(x),$ (2)
$h_{k}$ $=u_{k}+F_{k}(u_{k}),$ (3)
$u_{k+1}$ $=h_{k}+A_{k+1}(h_{k}),$ (4)
$h_{k+1}$ $=u_{k+1}+F_{k+1}(u_{k+1}),$ (5)
where $h_{k+1}=T_{\mathrm{seq}}(x)$ is the two-layer sequential output.
Layer Parallelism computation.
LP evaluates both attention modules at the shared input $x$, combines their outputs, and feeds the result to both FFN modules:
$\tilde{u}$ $=x+A_{k}(x)+A_{k+1}(x),$ (6)
[…]
This matches the (LP) equation in §3 and the implementation in Fig. 2(b), where divergent paths share intermediate states.
A.2 First-Order Error Analysis
The approximation error $\mathcal{E}(x)=T_{\mathrm{seq}}(x)-T_{\mathrm{LP}}(x)$ arises from evaluating submodules at different inputs. We decompose this error into three components.
Component 1: Attention evaluation shift.
In the sequential computation, $A_{k+1}$ is evaluated at $h_{k}=x+A_{k}(x)+F_{k}(u_{k})$, whereas LP evaluates it at $x$. Defining $\Delta_{1}=h_{k}-x=A_{k}(x)+F_{k}(u_{k})$, a first-order Taylor expansion yields:
$A_{k+1}(h_{k})-A_{k+1}(x)\approx J_{A_{k+1}}(x)\,\Delta_{1},$ (8)
where $J_{A_{k+1}}(x)$ is the Jacobian of $A_{k+1}$ at $x$.
Component 2: First FFN evaluation shift.
In the sequential computation, $F_{k}$ is evaluated at $u_{k}=x+A_{k}(x)$, whereas LP evaluates it at $\tilde{u}=x+A_{k}(x)+A_{k+1}(x)$. Since $\tilde{u}-u_{k}=A_{k+1}(x)$:
$F_{k}(u_{k})-F_{k}(\tilde{u})\approx-J_{F_{k}}(u_{k})\,A_{k+1}(x).$ (9)
Component 3: Second FFN evaluation shift.
In the sequential computation, $F_{k+1}$ is evaluated at $u_{k+1}=h_{k}+A_{k+1}(h_{k})$, whereas LP evaluates it at $\tilde{u}$. The difference is:
$u_{k+1}-\tilde{u}=F_{k}(u_{k})+\bigl[A_{k+1}(h_{k})-A_{k+1}(x)\bigr].$ (10)
Substituting the first-order approximation from (8):
$F_{k+1}(u_{k+1})-F_{k+1}(\tilde{u})\approx J_{F_{k+1}}(\tilde{u})\bigl[F_{k}(u_{k})+J_{A_{k+1}}(x)\,\Delta_{1}\bigr].$ (11)
Total error bound.
Combining equations (8)–(11) and taking norms:
$\boxed{\|\mathcal{E}(x)\|\lesssim\underbrace{\|J_{A_{k+1}}\|\,\|\Delta_{1}\|}_{\text{attention shift}}+\underbrace{\|J_{F_{k}}\|\,\|A_{k+1}(x)\|}_{\text{FFN}_{k}\text{ shift}}+\underbrace{\|J_{F_{k+1}}\|\bigl(\|F_{k}(u_{k})\|+\|J_{A_{k+1}}\|\,\|\Delta_{1}\|\bigr)}_{\text{FFN}_{k+1}\text{ shift}},}$ (12)
where Jacobian norms are operator norms evaluated at the appropriate inputs (suppressed for clarity).
A.3 Implications for Layer Selection
The error bound (12) reveals when LP introduces minimal degradation:
Small residual updates favor LP.
The bound depends on $\|\Delta_{1}\|=\|A_{k}(x)+F_{k}(u_{k})\|$, $\|A_{k+1}(x)\|$, and $\|F_{k}(u_{k})\|$. Layers where attention and FFN residuals are small relative to the residual stream contribute less error. This is consistent with the “residual stream dominance” phenomenon observed in deep transformers.
Low Jacobian sensitivity favors LP.
The terms $\|J_{A_{k+1}}\|$ and $\|J_{F_{k}}\|$, $\|J_{F_{k+1}}\|$ measure how sensitive each submodule is to input perturbations. Layers that are relatively insensitive to their exact input introduce less LP error.
Connection to CKA analysis (Fig. 5).
The CKA similarity between standard and counterfactual activations (with residual removed) serves as an empirical proxy for the Jacobian sensitivity. High CKA similarity indicates that the module output is relatively invariant to the upstream residual—precisely the condition under which the Jacobian terms in (12) are effectively small in the directions that matter. The plateau of high CKA values in mid-to-late layers (Fig. 5) corresponds to the region where LP is most effective (Fig. 4(e)).
Error accumulation through the network.
Errors injected at layer $k$ propagate through all subsequent layers via the Jacobian chain:
$\|\mathcal{E}_{\mathrm{output}}\|\lesssim\|\mathcal{E}_{k}\|\prod_{\ell>k}\|I+J_{f_{\ell}}\|,$ (13)
where $f_{\ell}$ is the full residual map of layer $\ell$. This implies:
  • Avoid early layers: Errors introduced early are amplified by many subsequent Jacobians.
  • Preserve a sequential tail: The final layers before the output logits are typically most sensitive (the language modeling head amplifies perturbations), so leaving them sequential stabilizes the output distribution.
A.4 Why Contiguous 2-Parallel Works Best
The “2-parallel” scheme (parallelizing consecutive pairs rather than arbitrary groups) succeeds because:
  1. 1.
    Local ordering matters less than global ordering. Shuffling experiments (Fig. 4(a)) show that adjacent layers are more interchangeable than distant ones, likely because they operate at similar levels of abstraction.
  2. 2.
    Error terms remain bounded. Within a single LP pair, the error is first-order in residual magnitudes. Chaining $n$ LP pairs gives $n$ independent first-order errors rather than a single large error from parallelizing all $2n$ layers simultaneously.
  3. 3.
    Intermediate synchronization corrects drift. Between LP pairs, the outputs are summed and re-normalized, preventing error accumulation within the LP region.
This explains why parallelizing triplets performs worse (as noted in §3): the second-order cross-terms become significant, and there is no intermediate synchronization to correct the trajectory.
A.5 Connection to GSM-8K Degradation
The disproportionate drop in GSM-8K accuracy under LP (Table 1) is consistent with recent findings that mathematical reasoning relies on sparse, localized circuits (stolfo2023mechanistic; christ2025mathneurosurgeryisolatinglanguage). These circuits likely have:
  • Larger effective Jacobians in the relevant directions (high sensitivity to precise intermediate states).
  • Less redundancy, so the LP approximation error is not absorbed by parallel pathways.
General language competence, by contrast, is distributed across many redundant pathways and is therefore more robust to the input perturbations introduced by LP.
Appendix B Comparison with Other Tensor Parallelism Optimizations
In this section, we compare Layer Parallelism (LP) with recent methods that aim to reduce communication overhead in tensor-parallel LLM inference: Sync-Point Drop (SPD) (kim2025spdsyncpointdropefficient), selective low-bit communication (dong2024lowbitcommunicationtensorparallel), and microscaling (MX) format compression (hansenpalmus2024communicationcompressiontensorparallel).
B.1 Method Overview
Sync-Point Drop (SPD).
kim2025spdsyncpointdropefficient selectively removes the all-reduce synchronization after the self-attention output projection, retaining only the FFN synchronization per block. They introduce modified block designs to minimize information loss and classify blocks into three sensitivity categories (in-sensitive, sensitive, extremely sensitive), applying block-to-block distillation to recover accuracy in sensitive layers.
Selective Low-bit Communication.
dong2024lowbitcommunicationtensorparallel compress communicated activations by quantizing most features to INT4 while keeping a small fraction ($1/64$) of high-range outlier features in BF16. This reduces communication from 16 bits to $\sim$4.2 bits per value on average, preserving outlier information critical for model performance.
MX Format Compression.
[…]
Layer Parallelism (Ours).
LP restructures the computational graph to execute consecutive layer pairs in parallel, reducing the number of sequential synchronization points. Unlike quantization-based methods, LP modifies the computation order rather than the communication encoding.
B.2 Quantitative Comparison
Table 3 compares results across methods. Direct comparison is challenging due to different models, hardware configurations, and evaluation metrics; we match the closest available configurations.
Table 3: Comparison of tensor parallelism optimization methods. Speedup is relative to standard tensor parallelism. For accuracy: SPD and LP report average zero-shot accuracy change; Low-bit reports performance retention; MX reports perplexity increase. Results marked are at 70% SPD application; indicates hardware-dependent results on PCIe-connected GPUs (L4); § indicates NVLink-connected GPUs (A100).
Method Model GPUs Speedup Accuracy Impact Mechanism
7B-scale Models
[…]
Low-bit LLaMA2-13B 8 99.5% retained Quant. comm
[…]
13B-scale Models
[…]
Low-bit LLaMA2-13B 8 99.5% retained Quant. comm
[…]
70B-scale Models
[…]
Table 4 provides a detailed accuracy comparison on common benchmarks where available.
Table 4: Accuracy comparison on zero-shot benchmarks. Values show absolute accuracy (%) or relative change from baseline. SPD results use ZS+B2B configuration at 70% SPD; Low-bit uses INT4+Selected BF16 at 4.2 bits.
Method Model ARC-e ARC-c HellaSwag WinoGrande Avg
Baseline (no optimization)
LLaMA2-13B 79.5 48.7 60.0 72.2 65.1
Gemma 2 27B 87.7 62.4 65.4 79.1 73.7
Low-bit Communication (dong2024lowbitcommunicationtensorparallel)
INT4+Sel. BF16 LLaMA2-13B 79.1 47.4 59.5 72.9 64.7
(-0.5%) (-2.8%) (-0.9%) (+1.0%) (-0.6%)
INT4+Sel. BF16 Gemma 2 27B 86.5 61.0 63.9 76.5 72.0
(-1.4%) (-2.2%) (-2.3%) (-3.3%) (-2.3%)
SPD (kim2025spdsyncpointdropefficient) at 70% blocks
ZS+B2B LLaMA2-13B (8-GPU) $\sim$65%
(<<1% drop)
ZS+B2B LLaMA2-70B (8-GPU) $\sim$66%
(-0.9%)
B.3 Key Observations
Complementary approaches.
[…]
Hardware sensitivity.
[…]
Model size scaling.
Both SPD and LP show improved robustness on larger models:
  • SPD: 44% of blocks are in-sensitive in LLaMA2-7B vs. 75% in LLaMA2-70B (kim2025spdsyncpointdropefficient)
  • LP: Larger models (Qwen3-14B) tolerate greater effective depth reduction than smaller models (LLaMA3.2-3B) before sharp accuracy drops (see Table 1)
Trade-off characteristics.
Each method exhibits distinct accuracy–speed trade-offs:
  • SPD: Graceful degradation; requires per-block sensitivity analysis and optional distillation
  • Low-bit: Minimal degradation ($\sim$0.5–2%) by preserving outlier features; no speedup measured
  • MX Compress: Perplexity increases 1–3%; speedup highly hardware-dependent
  • LP (Ours): Uniform accuracy drop across layers; GSM-8K disproportionately affected; fine-tuning recovers $\sim$50% of loss
B.4 Combining LP with Communication Compression
A promising direction is combining LP with quantization-based compression. Table 5 estimates potential combined benefits.
[…]
Optimization Sync Reduction Comm. Compression Est. Speedup
[…]
We leave empirical validation of combined LP + quantization to future work.
Appendix C Ablation: Tokens per second
@@
Appendix D Generalization to multiple GPUs
@@
Figure 10: Layer Parallelism in the case of parallelizing two layers over four accelerators. The stacked layers represent the tensor parallelism, and the colors indicate the processing of different previously contiguous layers. $Q,K,V,\text{att}\in\mathbb{R}^{T\times\frac{2D}{g}}$, where $D$ is the feature dimension and $g$ is the total number of accelerators.
@@
To confirm that the proposed scheme scales to commodity 4-GPU servers, we benchmark Llama 2 7B on a node with $4\times$ NVIDIA A100 80 GB PCIe accelerators while running the 4-GPU LP implementation. For each configuration we measure the wall-clock time of the decoding workload, normalize it to the vanilla $\Delta=32$ setting, and report the resulting relative throughput in Table 6. The gains steadily increase as we parallelize more layers (e.g., $\Delta=23$ reaches $1.46\times$), illustrating the benefit of halving the number of synchronization steps per block; however, these aggressive settings correspond to the larger accuracy drops discussed in the main text, so moderate $\Delta$ values offer a better accuracy–speed trade-off.
Table 6: Relative throughput of 4-GPU Layer Parallelism on Llama 2 7B measured on $4\times$ A100 80 GB PCIe GPUs. $\Delta$ denotes the number of layers replaced by LP pairs, with $\Delta=0$ serving as the baseline ($x1.00$).
$\Delta$ Rel. Speed
0 x1.00
10 x1.24
14 x1.34
18 x1.46
Appendix E Acceleration source
@@
Figure 11 illustrates flame graphs comparing two consecutive Llama 3.2 3B decoder layers using vanilla tensor parallelism and our LP approach. The profiling data summarized in Table 7 reveals that the primary source of acceleration in our LP method stems from reducing the total number of all-reduce synchronization operations across GPUs. Specifically, the vanilla tensor parallel approach performs synchronization at every decoder layer, resulting in higher cumulative synchronization overhead due to the fixed latency costs. Size-independent latency is dominated by GPU kernel-launch overhead, per-hop interconnect round-trip latency and host/stream synchronization imposed by the calling context. In contrast, our Layer Parallelism implementation runs pairs of layers simultaneously, effectively halving the number of synchronization points. This reduction in synchronization leads to a significant drop in synchronization time from 100.8ms to 50.7ms, directly contributing to the observed improvement in inference speed.